OpenStudy (anonymous):

Which of the following value pairs can be used as m and n to factor the expression 4x^2 - 8x + 3? A. 2, 6 B. –2, 6 C. 2, –6 D. –2, –6

hero (hero):

Hint: Find two numbers \(m\) and \(n\) so that: \(mn = 12\) \(m + n = -8\)

OpenStudy (anonymous):

I don't understand...

hero (hero):

Hint: Find two numbers m and n so that: \(m \times n = 12\) \(m + n = -8\) Whatever \(m\) and \(n\) are, they must satisfy both conditions. It's a matter of FINDING the two numbers. You ask yourself, "What two numbers can I think of that would multiply to get 12, but also add to get -8?

OpenStudy (anonymous):

A? (2, 6)

hero (hero):

2 + 6 = -8 ?

OpenStudy (anonymous):

not negative 8. anything I try to add with it just adds to -4...

OpenStudy (anonymous):

Wait, -2 and -6

hero (hero):

Yes -2 + -6 = -8 (-2)(-6) = 12 right?